# MatchU: Matching Unseen Objects for 6D Pose Estimation from RGB-D Images

### v1

@@ 1 Introduction
While recent works have extensively delved into generic pose estimation through learning models that rate and refine pose hypotheses [3, 33], their typical design involves time-consuming render-and-compare strategies. This constraint limits their utility in real-world applications. Another kind of method formulates the generic 6D pose estimation problem as a point cloud registration task, benefiting from the point cloud representation backbones but neglecting crucial texture information from RGB images [6, 68]. As a result, their descriptors learned solely from point clouds are usually less discriminative. Therefore, they introduce ambiguities in correspondence extraction partially remedied by adding knowledge about the object symmetry during training [68].
@@ 2 Related Work
The majority of related work focused on 6D pose estimation of seen objects for which training data (real or synthetic) is available. However, they need to be retrained for any new object instance. There were extension to object category pose estimation, but they cannot generalize to new unseen categories. Therefore, in recent two years several approaches, that aim at generalizing to novel unseen objects without retraining, were introduced.
@@
The approaches for seen object pose estimation rely on available real or synthetic training data and train one neural network model per object or per scene. They are usually multi-stage pipelines where the core learning efforts are in establishing image-to-model (2D-3D) correspondences further used for pose estimation through PnP+RANSAC or direct regression. A larger amount of learning based approaches consume RGB as input and only a few of them focus on RGB-D inputs facing the challenge of fusing RGB and depth information in neural networks.
RGB-D Fusion Methods are important because they profit from complementarity of two data sources and naturally improve pose accuracy as demonstrated in early works  [22, 65]. In deep learning approaches, features are extracted separately from two modalities with different neural networks and their fusion is not obvious. Early approach, like PointFusion [60], extracts global RGB (CNN) and depth (PointNet) features from the patch containing the object, and fuses them with per-point depth features for 3D object bounding box detection. Later, DenseFusion [54] performs late per-point feature fusion strengthened with the global information, allowing better discrimination at the local level and resulting in better occlusion handling. Other works like PVNet3D [19] rely on DenseFusion and estimate sparse keypoints instead of dense correspondences. FFB6D [21] instead uses bidirectional fusion modules to combine modality information at earlier stages and produce stronger per-pixel fused features. Recently, DFTr [69] uses Transformers and improves the data fusion with the global semantic similarity between RGB and depth. This fusion strategy can help handle missing and noisy data caused by reflections or low-texture information.
@@ 2.2 Unseen Object Pose Estimation
Template based methods like OSOP [47] and OVE6D [4] rely on representing the target object with templates. OSOP is a multi-stage pipeline, which leverages templates representing the 3D object seen from different views for segmentation, closest viewpoint selection and dense matching. OVE6D [4] is inspired by early ideas to learn specific embedding spaces for pose estimation [58], and, thus represents various 3D models together in an embedding space. ZePHyR [44] proposes hypothesis scoring, while MegaPose [33] proposes generic render-and-compare RGB pose refinement. The extracted features of OSOP [47] are not invariant to rotations and the design of the ZePHyR [44] and MegaPose [33] are computation-intensive since they need to evaluate many sampled hypotheses.
Natural way towards the generalization to unseen objects is through the descriptor learning, where generic descriptors can be used to match depth pixels of the object with its 3D model. Learning strong local 3D descriptors has been exhaustively studied in the context of point cloud registration [61, 26, 36, 45]. However, for object pose estimation descriptors need not only to be unique and repeatable, but also require rotation invariance introduced in RIGA [62], YOHO [57], and RoITr [64]. Moreover, adding color information to them and maintain generality is an additional challenge. Recent works tackling unseen object pose estimation, like Zeropose [6] and GCPose [68] rely on 3D descriptor learning. Zeropose [6] uses the foundation models ImageBind [15] and SAM [30] together with 3D-3D feature matching using descriptions from GeoTransformer [45]. GCPose [68] uses the same descriptor principle [45] together with explicit knowledge of object symmetries.
We leverage RoITr’s [64] rotation-invariance by design and do not require pre-defined object symmetries for training. We aim to get the best of both worlds by designing a rotation-invariant and symmetry aware backbone, which fuses RGB and depth information efficiently.
Figure 2: Overview of MatchU. Given an unseen object, we first extract its 3D and 2D local features from the CAD point cloud $P$, depth point cloud $Q$ as well as the RGB image $K$. The 3D and 2D local features are then fused in the latent space via our proposed Latent Fusion Attention Module with the supervision of a Bridged Coarse-level Matching Loss (BCM Loss)  $\mathcal{L}_{c}^{PKQ}$. The updated 3D descriptors $\widetilde{\phi}^{P^{\prime}}$ and $\widetilde{\phi}^{Q^{\prime}}$ are subsequently passed into the decoders that extract them into higher resolution of the descriptors ${\phi}^{P}$ and ${\phi}^{Q}$ with the supervision of a fine-level matching loss $\mathcal{L}^{PQ}_{f}$. Finally, the 6D pose of the unseen objects can be estimated by matching the descriptors in latent space and optimizing for the pose parameters $\mathcal{T}$.
@@ 3.1 Problem Formulation
$\mathcal{T^{*}}=\arg\min_{\mathcal{T}}\sum_{(p_{i},q_{j})\in\mathcal{C}}\|(\mathcal{T}p_{i}-q_{j})\|^{2}_{2}$ (1)
@@ 3.3 Encoding and Fusing Descriptors
Previous methods [21, 20, 54] usually interpolating features w.r.t. their spatial relationships explicitly. In MatchU, we use the positional encoding in the attention mechanism to incorporate the spatial awareness and implicitly align different modalities. For 2D features, we follow DETR [5] to encode the spatial information of the 2D feature map into the feature space. As for 3D, instead of encoding the raw position of the points [69], we propose to use the pose-agnostic Point Pair Features (PPFs) [13] as the position representation following  [63], which guarantees the geometric rotation-invariance and generalizability for unseen object pose estimation.
@@ 3.4 Learning to Describe
We adopt the Circle Loss [51], which maximizes the similarity of the positive pairs of the superpoints, as well as minimizes the similarity of the negative paris of the superpoints. Specifically, for each superpoint $p^{\prime}_{i}\in P^{\prime}$, and $q^{\prime}_{j}\in Q^{\prime}$, we can calculate the overlap $\mathcal{V}$ between $p^{\prime}_{i}$ and $q^{\prime}_{j}$ as:
@@
$\mathcal{L}_{c}^{PKQ}=(1-\lambda_{b})\mathcal{L}_{c}^{P^{\prime}Q^{\prime}}+\lambda_{b}(\mathcal{L}_{c}^{K^{\prime}Q^{\prime}}+\mathcal{L}_{c}^{K^{\prime}P^{\prime}}),$ (4)
where $\lambda_{b}$ is the weight for the 3D-3D and 3D-2D matching loss respectively.
@@
Table 1: Quantitative results 222 The numbers and timings are either from the original papers [6, 46] or BOP Challange [52]. in terms of Average Recall(AR) on BOP-5 core benchmark datasets for unseen object pose estimation task. MH: whether multiple hypotheses was adopted. Obj. Loc.: whether the object detector or segmentor are trained on the test objects. Refine.: whether the result is refined with either depth and/or RGB images.
@@ 4 Experiments and Results
We first evaluate our method on 5 BOP core datasets in comparison with other unseen object pose estimation methods in Section 4.2. Then in Section 4.3, we visualize our learned descriptors to highlight the capability of capturing the symmetry and encoding the texture simultaneously without external annotations. Since recent RGB-D fusion methods were designed only for seen objects, we compare our method with them on LM-O [2] dataset in Section 4.4. Ablation studies of the key designs of our method are conducted on LM-O dataset in Section 4.5. We introduce the implementation details and ablation studies of our verification as well as refinement in the Appendix.
@@ 4.1 Datasets and Evaluation Metrics
Figure 3: Qualitative results of 6D pose estimation of our method in comparison with Megapose and ZeroPose. The upper row shows an egg box which is heavily occluded in LM-O dataset. Our method is robust to handel occlusion while other methods flip the posese by mistake. The lower row shows an highly ambiguous object that other methods put the pose up side down but ours predicts accurately.
@@ 4.2 Evaluation on 5 BOP Core Benchmark Datasets
The evaluation results of our method are shown in Tab. 1. Compared with all the baselines, we achieve state-of-the-art performance on the task of unseen object pose estimation. We denote our method using only 1 pose hypothesis as Ours (Fast) and as Ours (Accurate) when multiple hypotheses are leveraged. For a fair comparison, we run our method under 5 specific settings in accordance with the baselines. In (a), we compare our method with ZeroPose [6], OSOP [47], and MegaPose [33] with only one hypothesis during the inference stage and without any refinement. Under (a), Ours (Fast) performs significantly better and faster on average. Note that in (a), we use a generic object detector for unseen objects objects [43]. This demonstrates the robustness of our method against noisy detection initialization, which often occurs in real applications. In (b), we further improve our results by introducing more pose hypotheses (20 by default) as we described in Section 3.5. We compare our method with OSOP [47] with the same number of hypotheses and exhibit superior performance. In (c), by adding the ICP refinement, our model achieves the best results among all the methods on both the overall and the per-dataset evaluation. Notably, in  (c), our method requires less time to perform one inference compared to other baselines [6, 33] with rendering-based refinement. In (d), we compare our method with OVE6D [4] and GCPose [68] with a seen-object detector. Our method consistently surpasses the baselines on T-LESS and TUD-L datasets. In (e), we finally improve our results by using a trained detector and incorporating ICP refinement. We showcase the qualitative results in Figure 3.
4.3 Capturing Symmetry and Texture
We visualize our descriptors for symmetric objects with t-SNE [53]. As shown in Figure  4, our learned descriptors can capture continuous (024_bowl) and discrete (#27_T-LESS) symmetries, which is credited to the rotation-invariant property of our design. Compared with GCPose [68], which relies on the supervision of symmetry labels, our descriptors recognize the symmetry even without any external symmetry annotations. Moreover, we visualize “tetX” from SYMSOL [42] dataset, a regular tetrahedron with one red and three white faces in Figure 5. The descriptors extracted by RoITr show the same distribution on all 4 faces due to their geometric similarity, introducing ambiguity issues for matching and leading to incorrect pose estimation potentially. In contrast, our method extracts distinct descriptors on the textured face. This indicates that our method not only describes the geometric property of the objects but also captures the texture information which further eliminates ambiguities in pose estimation.
@@ 4.4 Comparison with RGB-D Fusion Pipelines
4.5 Ablation Study
Key Design Principles.
@@ 4.5 Ablation Study
Influence of the Number of Hypotheses. As shown in Figure 6 (a), our method benefits from increasing the number of hypotheses. However, the performance saturates when the number passes 50, indicating 50 hypotheses could cover true poses for most cases. To balance the computation cost and performance, we use 20 as default.
@@ 5 Conclusion
We present MatchU, a Fuse-Describe-Match framework for unseen object pose estimation from single RGB-D images. Our method first extracts rotation-invariant descriptors from 3D point clouds of CAD model and depth map. Then, the multi-modal fusion of texture and geometry is achieved through a Latent Fusion Attention Module. A Bridged Coarse-Level Matching Loss is introduced to utilize latent features from RGB images to connect descriptions of partial observations and full object geometry. MatchU inherently captures object symmetry without explicit annotations. MatchU surpasses all existing methods for unseen object pose estimation by a large margin on standard benchmarks. Certainly, it relies on external object localization and could be could be affected by their erroneous results. In the future, incorporating such modules into the pipeline to build end-to-end training might further improve our results. We believe that by closing the gap to object-specific baselines, MatchU constitutes an important step forward to truly scalable 6D pose estimation of unseen objects.

### v2

@@ 1 Introduction
Recent works have extensively delved into generic pose estimation through learning models that rate and refine pose hypotheses [3, 30], predominantly utilizing time-consuming render-and-compare strategies. This constraint limits their utility in real-world applications. Alternatively, some approaches formulate the generic 6D pose estimation problem as a point cloud registration task, benefiting from the point cloud representation backbones but neglecting crucial texture information from RGB images [6, 65], leading to less distinct point cloud-based descriptors. Therefore, they introduce ambiguities in correspondence extraction partially remedied by adding knowledge about the object symmetry during training [65].
@@ 2 Related Work
The majority of related work focused on 6D pose estimation of seen objects for which training data (real or synthetic) is available. However, they need to be retrained for any new object instance. There were extensions to object category pose estimation, but they cannot generalize to new unseen categories. Therefore, in recent two years several approaches, that aim at generalizing to novel unseen objects without retraining, were introduced.
@@
The approaches for seen object pose estimation rely on available real or synthetic training data and train one neural network model per object or per scene. They are usually multi-stage pipelines where the core learning efforts are in establishing image-to-model (2D-3D) correspondences further used for pose estimation through PnP+RANSAC or direct regression. A larger amount of learning-based approaches consume RGB as input and only a few of them focus on RGB-D inputs facing the challenge of fusing RGB and depth information in neural networks.
RGB-D Fusion Methods are important because they profit from complementarity of two data sources and naturally improve pose accuracy as demonstrated in early works  [19, 62]. In deep learning approaches, features are extracted separately from two modalities with different neural networks and their fusion is not obvious. Early approache, like PointFusion [58], extracts global RGB (CNN) and depth (PointNet) features from the patch containing the object, and fuses them with per-point depth features for 3D object bounding box detection. Later, DenseFusion [52] performs late per-point feature fusion strengthened with the global information, allowing better discrimination at the local level and resulting in better occlusion handling. Other works like PVNet3D [17] rely on DenseFusion and estimate sparse keypoints instead of dense correspondences. FFB6D [18] instead uses bidirectional fusion modules to combine modality information at earlier stages and produce stronger per-pixel fused features. Recently, DFTr [66] uses Transformers and improves the data fusion with the global semantic similarity between RGB and depth. This fusion strategy can help handle missing and noisy data caused by reflections or low-texture information.
@@ 2.2 Unseen Object Pose Estimation
Template-based methods like OSOP [44] and OVE6D [4] rely on representing the target object with templates. OSOP is a multi-stage pipeline, which leverages templates representing the 3D object seen from different views for segmentation, closest viewpoint selection and dense matching. OVE6D [4] is inspired by early ideas to learn specific embedding spaces for pose estimation [56], and, thus represents various 3D models together in an embedding space. ZePHyR [42] proposes hypothesis scoring, while MegaPose [30] proposes generic render-and-compare RGB pose refinement. The extracted features of OSOP [44] are not invariant to rotations and the design of the ZePHyR [42] and MegaPose [30] are computation-intensive since they need to evaluate many sampled hypotheses.
A natural way towards the generalization to unseen objects is through descriptor learning, where generic descriptors can be used to match depth pixels of the object with its 3D model. Learning strong local 3D descriptors has been exhaustively studied in the context of point cloud registration [59, 23, 33, 43]. However, for object pose estimation descriptors need not only to be unique and repeatable, but also require rotation invariance introduced in RIGA [60], YOHO [55], and RoITr [61]. Moreover, adding color information to them and maintaining generality is an additional challenge. Recent works tackling unseen object pose estimation, like Zeropose [6] and GCPose [65] rely on 3D descriptor learning. Zeropose [6] uses the foundation models ImageBind [13] and SAM [27] together with 3D-3D feature matching using descriptions from GeoTransformer [43]. GCPose [65] uses the same descriptor principle [43] but with explicit knowledge of object symmetries.
We leverage RoITr’s [61] rotation-invariance by design and do not require pre-defined object symmetries for training. We aim to get the best of both worlds by designing a rotation-invariant and symmetry-aware backbone, which fuses RGB and depth information efficiently.
Figure 2: Overview of MatchU. Upon encountering an unseen object, we initially derive the segmented depth point cloud $Q$ and the corresponding RGB image crop $K$ utilizing a pre-trained generic segmentation network. Subsequently, we procure both 3D and 2D local features from the CAD point cloud $P$, depth point cloud $Q$, and the RGB image crop $K$. These extracted features are then amalgamated within a latent space through our innovative Latent Fusion Attention Module, under the guidance of a Bridged Coarse-level Matching Loss (BCM Loss) $\mathcal{L}{c}^{PKQ}$. The refined 3D descriptors $\widetilde{\phi}^{P^{\prime}}$ and $\widetilde{\phi}^{Q^{\prime}}$ are fed into decoders, which enhance the resolution of the descriptors to ${\phi}^{P}$ and ${\phi}^{Q}$, this process being steered by a detailed matching loss $\mathcal{L}^{PQ}_{f}$. In the final stage, the 6D pose of the novel objects is deduced by aligning the descriptors within the latent space and aggregating the pose parameters $\mathcal{T}$.
@@ 3.1 Problem Formulation
$\mathcal{T^{*}}=\arg\min_{\mathcal{T}\in\text{SE}(3)}\sum_{(p_{i},q_{j})\in\mathcal{C}}\|(\mathcal{T}p_{i}-q_{j})\|^{2}_{2}$ (1)
@@ 3.3 Encoding and Fusing Descriptors
Previous methods [18, 17, 52] usually interpolating features w.r.t. their spatial relationships explicitly. In MatchU, we use the positional encoding in the attention mechanism to incorporate spatial awareness and implicitly align different modalities. For 2D features, we follow DETR [5] to encode the spatial information of the 2D feature map into the feature space. As for 3D, instead of encoding the raw position of the points [66], we propose to use the pose-agnostic Point Pair Features (PPFs) [11] as the position representation following  [61], which guarantees the geometric rotation-invariance and generalizability for unseen object pose estimation.
@@ 3.4 Learning to Describe
We adopt the Circle Loss [48], which maximizes the similarity of the positive pairs of the superpoints, as well as minimizes the similarity of the negative pairs of the superpoints. Specifically, for each superpoint $p^{\prime}_{i}\in P^{\prime}$, and $q^{\prime}_{j}\in Q^{\prime}$, we can calculate the overlap $\mathcal{V}$ between $p^{\prime}_{i}$ and $q^{\prime}_{j}$ as:
@@
$\mathcal{L}_{c}^{PKQ}=\lambda_{b}\mathcal{L}_{c}^{P^{\prime}Q^{\prime}}+(1-\lambda_{b})(\mathcal{L}_{c}^{K^{\prime}Q^{\prime}}+\mathcal{L}_{c}^{K^{\prime}P^{\prime}}),$ (4)
where $\lambda_{b}$ and 1-$\lambda_{b}$ is the weight for the 3D-3D and 3D-2D matching loss respectively.
@@
Table 1: Quantitative results 222 The numbers and timings are either from the original papers [6, 44] or BOP Challange [50]. in terms of Average Recall(AR) on BOP-5 core benchmark datasets for unseen object pose estimation task. MH: whether multiple hypotheses were adopted. Obj. Loc.: whether the object localization (detection or segmentation) is trained on the test objects. Refine.: whether the result is refined with either depth and/or RGB images.
@@ 4 Experiments and Results
4.1 Implementation Details
During training, default settings are $\lambda_{b}=0.3$ and $\lambda_{c}=0.5$. For inference, we set $\kappa=128$, $s=64$, and initially $\eta=20$, adjusting $\eta$ to 64 to enhance accuracy during testing. To localize unseen objects in RGB-D images, we utilize a CAD-based segmentation method [41]. For 6D pose evaluation with instance-level localization, existing detection results from [6, 34, 49] are employed. Detailed implementation and more information about the network designs are provided in the Appendix.
@@ 4.3 Evaluation on 5 BOP Core Benchmark Datasets
The evaluation results of our method are shown in Tab. 1. Compared with all the baselines, we achieve state-of-the-art performance on the task of unseen object pose estimation. We denote our method using only 1 pose hypothesis as Ours (Fast) and as Ours (Accurate) when multiple hypotheses are leveraged. For a fair comparison, we run our method under 5 specific settings in accordance with the baselines. In (a), we compare our method with ZeroPose [6], OSOP [44], and MegaPose [30] with only one hypothesis during the inference stage and without any refinement. Under (a), Ours (Fast) performs significantly better and faster on average. Note that in (a), we use a generic instance detection/segmentation network for unseen objects [41]. This demonstrates the robustness of our method against noisy detection initialization, which often occurs in real applications. In (b), we further improve our results by introducing more pose hypotheses (20 by default) as we described in Section 3.5. We compare our method with OSOP [44] with the same number of hypotheses and exhibit superior performance. In (c), by adding the ICP refinement, our model achieves the best results among all the methods on both the overall and the per-dataset evaluation. Notably, in  (c), our method requires less time to perform one inference compared to other baselines [6, 30] with rendering-based refinement. In (d), we compare our method with OVE6D [4] and GCPose [65] with an identical object detection/segmentation network. Our method consistently surpasses the baselines on T-LESS and TUD-L datasets. In (e), we finally improve our results by using a trained detector and incorporating ICP refinement. We showcase the qualitative results in Figure 5.
@@
4.4 Capturing Symmetry and Texture
We visualize our descriptors for symmetric objects with t-SNE [51]. As shown in Figure  3, our learned descriptors can capture continuous (024_bowl) and discrete (#27_T-LESS) symmetries, which is credited to the rotation-invariant property of our design. Compared with GCPose [65], which relies on the supervision of symmetry labels, our descriptors recognize the symmetry even without any external symmetry annotations. Moreover, we visualize “tetX” from SYMSOL [40] dataset, a regular tetrahedron with one red and three white faces in Figure 4. The descriptors extracted by RoITr show the same distribution on all 4 faces due to their geometric similarity, introducing ambiguity issues for matching and leading to incorrect pose estimation potentially. In contrast, our method extracts distinct descriptors on the textured face. This indicates that our method not only describes the geometric property of the objects but also captures the texture information which further eliminates ambiguities in pose estimation.
@@ 4.5 Comparison with RGB-D Fusion Pipelines
Figure 5: Qualitative results of 6D pose estimation of our method in comparison with Megapose and ZeroPose. The upper row shows an egg box which is heavily occluded in the LM-O dataset. Our method is robust to handle occlusion while other methods flip the poses by mistake. The lower row shows a highly ambiguous object that other methods put the pose upside down but ours predicts accurately.
4.6 Ablation Study
Key Design Principles.
@@ 4.6 Ablation Study
Influence of the Number of Hypotheses. As shown in Figure 6 (a), our method benefits from increasing the number of hypotheses. However, the performance saturates when the number passes 50, indicating 50 hypotheses could cover true poses for most cases. To balance the computation cost and performance, we use 20 as the default.
@@ 5 Conclusion
We present MatchU, a Fuse-Describe-Match framework for unseen object pose estimation from single RGB-D images. Our method first extracts rotation-invariant descriptors from 3D point clouds of CAD model and depth map. Then, the multi-modal fusion of texture and geometry is achieved through a Latent Fusion Attention Module. A Bridged Coarse-Level Matching Loss is introduced to utilize latent features from RGB images to connect descriptions of partial observations and full object geometry. MatchU inherently models object symmetry without explicit annotations. MatchU surpasses all existing methods for unseen object pose estimation by a large margin on standard benchmarks. Certainly, it relies on external object localization and could be could be affected by their erroneous results. In the future, incorporating such modules into the pipeline to build end-to-end training might further improve our results. We believe that by closing the gap to object-specific baselines, MatchU constitutes an important step forward to truly scalable 6D pose estimation of unseen objects.